OpenStudy (unikornkiller):

A water sprinkler sends water out in a circular pattern. How large is the watered area if the radius of the watering pattern is 19 ft?

OpenStudy (unikornkiller):

its 23 ft

Directrix (directrix):

The area of the watering circle is pi * r^2 =pi * 19^2. What is that equal to: pi * 19 * 19 =

Directrix (directrix):

@UniKorNKiLlEr